### v1

# Breaking the Orthogonality Barrier in Quantum LDPC Codes

Abstract
Classical low-density parity-check (LDPC) codes are a widely deployed and well-established technology, forming the backbone of modern communication and storage systems. It is well known that, in this classical setting, increasing the girth of the Tanner graph while maintaining regular degree distributions leads simultaneously to good belief-propagation (BP) decoding performance and large minimum distance. In the quantum setting, however, this principle does not directly apply because quantum LDPC codes must satisfy additional orthogonality constraints between their parity-check matrices. When one enforces both orthogonality and regularity in a straightforward manner, the girth is typically reduced and the minimum distance becomes structurally upper bounded.
In this work, we overcome this limitation by using permutation matrices with controlled commutativity and by restricting the orthogonality constraints to only the necessary parts of the construction, while preserving regular check-matrix structures. This design breaks the conventional trade-off between orthogonality, regularity, girth, and minimum distance, allowing us to construct quantum LDPC codes with large girth and without the usual distance upper bounds. As a concrete demonstration, we construct a girth-8, (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$.
1 Introduction
Classical low-density parity-check (LDPC) codes are a widely deployed and well-established technology in modern communication and storage systems [7], and BP decoding can achieve channel capacity [16, 25]. Performance depends strongly on the Tanner-graph degree distribution: regular codes can perform well, and carefully designed irregular distributions can further improve BP performance [24], whereas naive irregularization can degrade it. Random LDPC codes with variable-node degree at least 3 have minimum distance growing linearly with blocklength [26, 7], so classical LDPC research has not focused on increasing minimum distance itself. Moreover, BP decoding can be trapped by small Tanner-graph structures called trapping sets [27], which cause decoding failures; suppressing harmful trapping sets typically requires large girth [29].
Sparse-graph quantum error-correcting codes were proposed in [18], and the behavior of iterative decoding and the role of degeneracy were discussed in [23]. Families with positive rate and distance $\Theta(\sqrt{n})$ are known [31]. Constructions based on Kronecker sum and product [15], distance bounds for generalized bicycle (GB) codes [32], finite-length evaluations [20], and decoding protocols [17] have been reported, advancing the theoretical understanding of CSS-type QLDPC codes. However, the orthogonality constraint between the CSS parity-check matrices $H_{X}$ and $H_{Z}$, namely $H_{X}H_{Z}^{\mathsf{T}}=0$, has no classical counterpart, and enforcing both orthogonality and regularity in a straightforward manner typically reduces girth and induces structural distance upper bounds.
Surface codes and hypergraph product (HGP) codes rely on geometric or product structures and thus follow design principles different from those developed for classical LDPC codes. As a result, it is not straightforward to apply classical degree-distribution and girth design techniques directly [6, 31, 26]. The goal of this work is to provide a construction principle that preserves classical LDPC parity-check structures while enabling their direct use as CSS check matrices. Specifically, we aim for regular LDPC codes with large girth and variable degree at least 3, whose minimum distance is not trivially upper bounded.
Prior work has explored cyclic parity-check matrices that allow explicit control of regular Tanner-graph degree distributions. Representative constructions include the Hagiwara–Imai codes [10] and bicycle constructions [18, 15, 32]. However, these constructions have known limitations. Even for classical codes, one-step lifting of a protograph (lifting by circulant permutation matrices (CPMs)) can impose fixed upper bounds on girth due to the protograph structure. Mitchell et al. pointed out that one-step lifting can impose fixed bounds on minimum distance and girth and that two-step lifting can improve them [19]. For CSS codes, the parity-check matrices $H_{X}$ and $H_{Z}$ are QC, and the Tanner-graph girth satisfies $g=\min\{g(H_{X}),g(H_{Z})\}$, so QC girth bounds apply to each matrix. For quasi-cyclic CSS designs, generalized CSS (entanglement-assisted) constructions use classical QC constructions with girth $\geq 6$ to avoid 4-cycles [11]. Explicit constructions achieving girth 12 for column weight 2 have also been reported [14]. Moreover, for general quantum CPM–LDPC codes, the girth is upper bounded by 12 for column weight 2 and by 6 for column weight at least 3 [1]. Using APM–LDPC codes, these upper bounds can be broken [14, 13].
In CSS-type LDPC constructions, deleting rows from the parity-check matrices $H_{X}$ or $H_{Z}$ preserves the commutation condition $H_{X}H_{Z}^{\mathsf{T}}=0$, but it weakens stabilizer constraints and can enlarge the code space [18, 10]. Row deletion is used for rate adjustment and for controlling row and column weights, yet distance preservation is not guaranteed in general [10, 22]. In particular, deleting rows can allow low-weight vectors to enter $C_{X}\setminus C_{Z}^{\perp}$ or $C_{Z}\setminus C_{X}^{\perp}$, creating new low-weight logical operators and thus reducing the minimum distance. In CSS codes, with $C_{X}=\ker(H_{X})$ and $C_{Z}=\ker(H_{Z})$, low-weight vectors in $C_{X}\setminus C_{Z}^{\perp}$ or $C_{Z}\setminus C_{X}^{\perp}$ become non-trivial logical operators [3, 28, 9]. Consequently, deleted check rows (which are LDPC and thus low weight) can become logical operators and upper bound the distance by the row weight. In the CSS product-code constructions of Ostrev et al., row removal is explicitly used to control the number of stabilizers, and distance degradation can occur [22]. Lin et al. report that removing redundant stabilizer rows reduces the syndrome distance, which is a structural effect due to fewer constraints, not a decoder artifact [17].
In this work, we first discuss the general mechanism by which row deletion can degrade the minimum distance of CSS codes. Next, for generalized Hagiwara–Imai codes [14], we establish a code-design method that controls commutativity of submatrices to avoid distance degradation due to row deletion. More concretely, using the algebraic structure of APM–LDPC codes, we guarantee orthogonality only on the active part, introduce a set of row pairs for which commutation is required, and localize the commutation condition. We further control the nonzero pattern of the interaction matrix so that deleted rows do not materialize as low-weight logical operators, and present a sequential construction algorithm based on the APM composition rule. As a concrete demonstration, we construct a (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate (FER) as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$. We also compare constructions with and without commutation control to demonstrate suppression of low-weight logical operators.
2 Problem setting: low-weight logical operators induced by check-row removal
We abstract the conventional CSS–LDPC construction used in this paper, based on [18, 10]. We first construct two orthogonal square (block-circulant) mother matrices $\hat{H}_{X},\hat{H}_{Z}$. The CSS code defined by these mother matrices typically has rate zero, so to adjust the rate we delete internal rows and use the remaining rows as the active matrices $H_{X}$ and $H_{Z}$. Namely,
$\hat{H}_{X}=\begin{bmatrix}H_{X}\\ \tilde{H}_{X}\end{bmatrix},\qquad\hat{H}_{Z}=\begin{bmatrix}H_{Z}\\ \tilde{H}_{Z}\end{bmatrix}$ (1)
We refer to $H_{X}$ and $H_{Z}$ as the active part and $\tilde{H}_{X}$ and $\tilde{H}_{Z}$ as the latent part.
However, in this construction, mother orthogonality imposes strong constraints on the latent rows and can degrade the minimum distance. The minimum distance is defined as $d_{\min}=\min\{d_{X},d_{Z}\}$ with
$d_{Z}:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in C_{X}\setminus C_{Z}^{\perp}\},\qquad d_{X}:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in C_{Z}\setminus C_{X}^{\perp}\}.$ (2)
Mother orthogonality
$\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}=0$ (3)
forces not only the active orthogonality between the active matrices $H_{X}$ and $H_{Z}$, namely $H_{X}H_{Z}^{\mathsf{T}}=0$, but also
$H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=0,\qquad H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=0$ (4)
which implies $\operatorname{Row}(\tilde{H}_{X})\subset C_{Z}$ and $\operatorname{Row}(\tilde{H}_{Z})\subset C_{X}$. Thus each row $\boldsymbol{x}$ of $\tilde{H}_{X}$ satisfies $\boldsymbol{x}\in C_{Z}$, and each row $\boldsymbol{z}$ of $\tilde{H}_{Z}$ satisfies $\boldsymbol{z}\in C_{X}$. In general $\boldsymbol{x}$ need not belong to $C_{X}^{\perp}$ and $\boldsymbol{z}$ need not belong to $C_{Z}^{\perp}$, in which case they become logical operators. The minimum row weight (in our construction, $L$) becomes an upper bound on $d_{X}$ and $d_{Z}$.
In this work we also obtain the active matrices $H_{X},H_{Z}$ by deleting rows from the mother matrices, but we design them so that low-weight rows in the latent matrices are not orthogonal to $H_{Z}$ or $H_{X}$, respectively. That is, for low-weight $\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})$ we enforce $H_{Z}\boldsymbol{x}^{\mathsf{T}}\neq 0$, and for low-weight $\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})$ we enforce $H_{X}\boldsymbol{z}^{\mathsf{T}}\neq 0$, so that in particular $\operatorname{Row}(\tilde{H}_{X})\not\subset C_{Z}$ and $\operatorname{Row}(\tilde{H}_{Z})\not\subset C_{X}$. The same argument applies not only to individual latent rows but also to low-weight linear combinations.
Accordingly, we define latent-based upper bounds for the $X$ and $Z$ distances as
$d_{X}^{(\mathrm{res})}$ $:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})\cap C_{Z}\setminus C_{X}^{\perp}\},$ (5)
$=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in C_{Z},\ \boldsymbol{x}=(\tilde{H}_{X})^{\mathsf{T}}\boldsymbol{u},\ \text{for some }\boldsymbol{u},\ \boldsymbol{x}\notin C_{X}^{\perp}\},$ (6)
$d_{Z}^{(\mathrm{res})}$ $:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})\cap C_{X}\setminus C_{Z}^{\perp}\},$ (7)
$=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in C_{X},\ \boldsymbol{z}=(\tilde{H}_{Z})^{\mathsf{T}}\boldsymbol{v},\ \text{for some }\boldsymbol{v},\ \boldsymbol{z}\notin C_{Z}^{\perp}\}$ (8)
The proposed design aims to make these bounds as large as possible. In [12], binary submatrices corresponding to codewords are lifted to a nonbinary field so that they no longer represent codewords; this can be viewed as increasing these residual-based bounds via nonbinary lifting.
3 Method of matrix design
In this section, based on the above problem setting, we aim to construct the latent matrices $\tilde{H}_{X}$ and $\tilde{H}_{Z}$ such that
$H_{X}H_{Z}^{\mathsf{T}}=0,\qquad H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\neq 0,\qquad H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\neq 0$ (9)
and $d_{X}^{(\mathrm{res})},d_{Z}^{(\mathrm{res})}$ are large. We define generalized Hagiwara–Imai codes, i.e., mother matrix pairs with block-circulant structure, and develop a general theory for imposing only the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ between the active matrices $H_{X}$ and $H_{Z}$.
A protograph LDPC code with column weight $J$ and row weight $L$ is defined by a parity-check matrix consisting of $J\times L$ permutation matrices of size $P$ [30]. For a permutation $f:[P]\to[P]$, the corresponding $P\times P$ permutation matrix $F=P(f)$ is defined by $F_{x,y}=1\;\Longleftrightarrow\;f(x)=y.$ Each row and column has exactly one 1, so the mother matrix is sparse while its block structure is compact.
We introduce generalized Hagiwara–Imai codes [14] as protograph generalizations of the Hagiwara–Imai quasi-cyclic CSS codes [10]. For permutation matrices $F_{i},G_{i}\ (i\in[L_{2}])$ of size $P$, define the mother matrices $\hat{H}_{X},\hat{H}_{Z}$ of size $(L_{2}P)\times(LP)$ by
$(\hat{H}_{X})_{i,j}=F_{j-i},\quad(\hat{H}_{X})_{i,L_{2}+j}=G_{j-i},$ (10)
$(\hat{H}_{Z})_{i,j}=(G_{i-j})^{\mathsf{T}},\quad(\hat{H}_{Z})_{i,L_{2}+j}=(F_{i-j})^{\mathsf{T}}.$ (11)
All subscripts are taken modulo $L_{2}$. Throughout, indices in $[L/2]$ are taken modulo $L/2$, i.e., we identify $[L/2]$ with $\mathbb{Z}_{L/2}$. A sufficient condition for orthogonality is given below. The active matrices $H_{X}$ and $H_{Z}$ consist of the top $J$ block rows. Conventional designs impose commutativity of $F_{i}$ and $G_{j}$ so that both $H_{X},H_{Z}$ and $\hat{H}_{X},\hat{H}_{Z}$ are orthogonal. The top $J$ block rows of the mother matrices are used as the active matrices $H_{X}$ and $H_{Z}$. Our interest is to satisfy active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ while not imposing mother-matrix orthogonality.
Each block row of $\hat{H}_{X}$ is a cyclic shift of $(F_{0},F_{1},\dots,F_{L/2-1})$ in the left half and $(G_{0},G_{1},\dots,G_{L/2-1})$ in the right half. Hence blocks depend only on the difference (ji)(j-i), and the product of mother matrices depends only on the difference. For any $i,k\in[L_{2}]$, the $(i,k)$ block is
$\bigl(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}\bigr)_{i,k}=\sum_{u=0}^{L_{2}-1}\Bigl(F_{u}G_{k-u}+G_{k-u}F_{u}\Bigr)=:\Psi_{r}$ (12)
which depends only on $r=(k-i)\bmod\,L_{2}$. For $r\in[L/2]$, if $F_{u}$ and $G_{r-u}$ commute for all $u\in[L/2]$, then $\Psi_{r}=0$.
3.1 Sufficient condition for active orthogonality
We re-derive a convenient form of a sufficient condition for $H_{X}H_{Z}^{\mathsf{T}}=0$ [12]. Define
$\Delta:=\{(k-i)\bmod\,L/2\mid 0\leq i,k\leq J-1\}\subseteq[L/2].$ (13)
The next theorem spells out how commutativity restricted to these differences guarantees active orthogonality.
Theorem 3.1.
If $F_{u}$ and $G_{r-u}$ commute for all $r\in\Delta$ and all $u\in[L/2]$, then $H_{X}H_{Z}^{\mathsf{T}}=0$.
Proof.
For any $i,k\in[J]$, we have $r=(k-i)\bmod\,L/2\in\Delta$. By assumption, $\Psi_{r}=0$, hence for all $i,k\in[J]$, $(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}})_{i,k}=\Psi_{(k-i)\bmod\,L/2}=0.$ Therefore $H_{X}H_{Z}^{\mathsf{T}}=0$. ∎
3.2 Necessary condition for latent non-orthogonality
Define the index pairs of $(F_{i},G_{j})$ required to commute by
$\Gamma:=\bigcup_{r\in\Delta}\Gamma_{r},\qquad\Gamma_{r}:=\{(i,j)\mid(i,j)=(u,r-u),\ u\in[L/2]\}$
where subscripts are modulo $L/2$. Here $\Gamma_{r}$ is the set of index pairs contributing to $\Psi_{r}$, and $\Gamma_{r}\cap\Gamma_{s}=\varnothing$ for $r\neq s$.
If $J\geq L/2$ then the rate is zero, so we assume $J<L/2$. We first state a basic feasibility constraint: without enough blocks, latent non-orthogonality cannot be forced.
Theorem 3.2.
Assume the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$. For the standard active choice (top $J$ block rows), to have $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\neq 0$ and $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\neq 0$, it is necessary that $L\geq 4J$.
Proof.
Assume $L<4J$. Then for this active choice, $\Delta=[L/2]$. Active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ implies $\Psi_{r}=0$ for all $r\in\Delta$, hence $\Psi_{r}=0$ for all $r\in[L/2]$. Therefore every block of $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ vanishes, in particular the mixed blocks between active and latent rows, so $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=0$ and $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=0$, contradicting the premise. ∎
We next isolate the minimal algebraic obstruction to latent non-orthogonality.
Theorem 3.3.
To have $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\neq 0$ and $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\neq 0$, it is necessary that there exists $r\in[L/2]\setminus\Delta$ with $\Psi_{r}\neq 0$.
Proof.
Assume $\Psi_{r}=0$ for all $r\in[L/2]\setminus\Delta$. For any $i\in[L_{2}]\setminus[J]$ and $k\in[J]$, $r=(k-i)\bmod\,L_{2}$ does not belong to $\Delta$. By assumption $\Psi_{r}=0$, hence for all such $i,k$, $(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}})_{i,k}=\Psi_{(k-i)\bmod\,L_{2}}=0.$ Therefore $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=0$, and similarly $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=0$, a contradiction. If all pairs $(i,j)\in[L/2]^{2}\setminus\Delta$ commute, then each term $F_{u}G_{r-u}+G_{r-u}F_{u}$ vanishes over $\mathbb{F}_{2}$, so $\Psi_{r}=0$ for all $r\notin\Delta$. ∎
3.3 Upper bound on girth
Finally, we recall a structural limitation that applies when commutativity holds on all required differences.
Theorem 3.4.
Let $L$ be even and assume $2\leq J\leq L/2$. If $F_{u}$ and $G_{r-u}$ commute for all $r\in\Delta$ and all $u\in[L/2]$, then the Tanner graphs of $H_{X}$ and $H_{Z}$ each contain an 8-cycle.
Proof.
Since $J\geq 2$, we have $0,1\in\Delta$. Also $L/2\geq 2$, so there exist $i\neq i^{\prime}$ in $[L/2]$. Let $r=0$, $s=1$, and set j=rij=r-i, $j^{\prime}=s-i$ (indices modulo $L/2$). By definition, $(i,j),(i,j^{\prime})\in\Gamma_{r}\cup\Gamma_{s}$, and similarly $(i^{\prime},j),(i^{\prime},j^{\prime})\in\Gamma_{r}\cup\Gamma_{s}$. By assumption the corresponding $F$ and $G$ commute, so the 8-cycle word composed of the two $F$ columns and two $G$ columns is
$W=F_{i}G_{j}^{-1}G_{j^{\prime}}F_{i}^{-1}F_{i^{\prime}}G_{j^{\prime}}^{-1}G_{j}F_{i^{\prime}}^{-1}$ (14)
Using commutativity to reorder gives $W=(F_{i}F_{i}^{-1})(F_{i^{\prime}}F_{i^{\prime}}^{-1})(G_{j}^{-1}G_{j^{\prime}})(G_{j^{\prime}}^{-1}G_{j})=I$, so an 8-cycle necessarily exists in the Tanner graphs of $H_{X}$ and $H_{Z}$. ∎
4 Method of Matrix Construction
Building on the commutation conditions and active orthogonality derived in the previous section, we now present a concrete sequential procedure to construct permutation blocks that satisfy those constraints while avoiding short cycles. For a given $(L/2,P,J)$, we construct $\{F_{i}\}$ and $\{G_{i}\}$ that satisfy the following simultaneously:
  • (1)
    $F_{i}$ and $G_{j}$ commute for $(i,j)\in\Delta$.
  • (2)
    At least one $(i,j)\in[L/2]^{2}\setminus\Delta$ is non-commuting.
  • (3)
    Avoid short cycles in the active matrices $H_{X}$ and $H_{Z}$.
The framework of constructing quasi-cyclic LDPC blocks from circulant permutation matrices (CPMs) was systematized by Fossorier [5]. We adopt affine permutation matrices (APMs): the APM–LDPC framework extends CPMs, and its algebraic form makes commutativity control straightforward via congruence conditions [8]. In the classical setting, Yoshida and Kasai reported that linear permutation polynomial (APM-based) codes achieve performance comparable to protograph LDPC codes [33], while no fixed girth upper bound has been reported for APM-based constructions, unlike one-step CPM lifting. Methods that combine APM–LDPC codes to extend length and girth have also been proposed [21]. From a group-theoretic viewpoint, APMs on $\mathbb{Z}_{P}$ form the affine group $\mathrm{AGL}(1,\mathbb{Z}_{P})=\mathbb{Z}_{P}\rtimes\mathbb{Z}_{P}^{\times}$, a semidirect product of translations by the unit group; commutativity is therefore governed by the interaction between the linear and translation parts.
Consider an affine permutation (AP) on $\mathbb{Z}_{P}$,
$f(x)=ax+b,\qquad a\in\mathbb{Z}_{P}^{\times},\ b\in\mathbb{Z}_{P}.$
We set $f_{i}(x)=a_{i}x+b_{i}$ and $g_{j}(x)=c_{j}x+d_{j}$, and denote the corresponding permutation matrices by $F_{i}:=P(f_{i})$ and $G_{j}:=P(g_{j})$.
Under this representation, commutativity of APs can be checked by a quadratic congruence:
$f_{i}g_{j}=g_{j}f_{i}\iff d_{j}(a_{i}-1)-b_{i}(c_{j}-1)\equiv 0\pmod{P}$, (C(i,j))
See [8] for a derivation. The condition involves products of $a_{i},c_{j}$ and $b_{i},d_{j}$ [8]. The commutation table can thus be expressed as a system of congruences; in particular, if $a_{i},c_{j}$ are chosen first, the commutation constraints reduce to linear congruences in $b_{i},d_{j}$, enabling a consistent search [8, 21].
We sequentially select candidates that satisfy both the commutation table and short-cycle conditions, using backtracking when a candidate fails. The number of trials is adjusted dynamically based on recent success rates. We interpret each candidate generator as an arm in a multi-armed bandit and allocate trials accordingly [2].
Short-cycle detection reduces to checking fixed points of the composite map $\Sigma$ along a block-cycle pattern [5, 8]. The AP composition is again AP, $\Sigma(x)=Ax+B$, so
$\Sigma(x)=x\ \Longleftrightarrow\ \gcd(A-1,P)\mid B$
provides a test without enumerating the Tanner graph [8].
5 Method of Decoding
Our experiments use BP decoding [18, 23] as the baseline. We first recall a basic principle: if we knew the support where the BP estimate differs from the true error, decoding would reduce to a linear system on the corresponding columns, and a unique solution would imply successful recovery. For example, if the $X$-side mismatch support is $E$, then $H_{Z}(:,E)\,\boldsymbol{e}_{E}=\boldsymbol{r}_{X}$; if $H_{Z}(:,E)$ has full column rank, the solution is unique and yields the correct correction. The same applies to the $Z$ side with $H_{X}$ and $\boldsymbol{r}_{Z}$.
Let the BP estimate at each iteration be $\boldsymbol{\hat{x}},\boldsymbol{\hat{z}}$. Define the residual syndromes
$\boldsymbol{r}_{X}:=\boldsymbol{s}_{X}\oplus H_{Z}\boldsymbol{\hat{x}},\qquad\boldsymbol{r}_{Z}:=\boldsymbol{s}_{Z}\oplus H_{X}\boldsymbol{\hat{z}}$.
Empirically, BP stalls or slowly oscillates roughly once per $10^{5}$ trials, and we trigger post-processing in such cases. We invoke three post-processing methods: ETS-based post-processing, a correction based on flip history, and an OSD-based method. Below we briefly describe the procedure and the criteria. In post-processing, the residual syndrome $\boldsymbol{r}$ is $\boldsymbol{r}_{X}$ (with $H=H_{Z}$) or $\boldsymbol{r}_{Z}$ (with $H=H_{X}$), depending on the side.
5.1 Flip history decoding
Flip history decoding (FHD) forms a candidate set $F$ from variables that flipped during BP and seeks a correction that cancels the residual syndrome on the local submatrix. This local correction is a heuristic to resolve stalls caused by trapping structures [27]. Related local-flip decoders include weighted bit-flipping [34]. Concretely, let $N(F)$ be the set of check rows adjacent to $F$, and solve the GF(2) linear system
$H_{N(F),F}\,\boldsymbol{\delta}=\boldsymbol{r}_{N(F)}.$
If the system is solvable and the solution is unique, we flip the bits indicated by $\boldsymbol{\delta}$.
5.2 OSD
OSD [4] uses the least reliable variables from BP to form a column set $K$ and solves
$H_{:,K}\,\boldsymbol{x}=\boldsymbol{r}$
using all rows. The method is applied independently to the $X$ and $Z$ sides. In the implementation, we perform a binary search on $|K|$ to find the minimum $K$ that is solvable, then take the largest $K$ within the range that preserves uniqueness. The correction is applied only when the solution weight is below a threshold.
5.3 Elementary Trapping Set (ETS)-based post-processing
ETS-based post-processing targets elementary trapping sets with $b=2$ unsatisfied checks, which are among the most harmful structures that cause BP to stall. We precompute a library of ETSs from the Tanner graph (e.g., $(6,2)$, $(8,2)$, $(10,2)$) and store each entry as a variable set $V$ and its odd check pair $(c_{0},c_{1})$. During decoding, we run ETS-based post-processing only when the number of unsatisfied checks is small (we skip post-processing when the number is not small) and the residual syndrome has exactly two unsatisfied checks.
Given the current estimate, let $\boldsymbol{r}$ be the residual syndrome and $U=\{c_{0},c_{1}\}$ be its unsatisfied check pair. We scan ETS entries whose odd checks match $U$. For each candidate ETS with variable set $V$, we build the local check set $N(V)$ and reject it if $\boldsymbol{r}$ has nonzero entries in $N(V)\setminus U$. Otherwise we solve the local system
$H_{N(V),V}\,\boldsymbol{\delta}=\boldsymbol{r}_{N(V)}$
over $\mathbb{F}_{2}$ using Gaussian elimination. If solvable, we flip variables in $V$ according to $\boldsymbol{\delta}$ and accept the correction only when the syndrome matches the target. We try ETS sizes in increasing order and apply the first successful correction, independently for the $X$ and $Z$ sides.
6 Results
We instantiate the general theory of Section 3 and the sequential construction of Section 4 for the smallest case we were able to construct, $J=3,L=12,P=768$. We first give explicit $\Delta,\Gamma$ and $\Psi_{r}$, clarify the active orthogonality condition for the active matrices $H_{X}$ and $H_{Z}$, then design the commutation table and APM parameters, and finally evaluate distance bounds and FER.
6.1 Structure and active orthogonality
For $J=3,L/2=6$, each block row is a cyclic shift of the previous row, so $\hat{H}_{X},\hat{H}_{Z}$ have a $6\times 12$ block-circulant structure:
$\begin{aligned} \hat{H}_{X}=&\left(\begin{array}[]{llllll|llllll}F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}\\ F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}\\ F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}\\ \hline\cr F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}\\ F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}\\ F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}\end{array}\right).\\ \hat{H}_{Z}=&\left(\begin{array}[]{llllll|llllll}G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}\\ G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}\\ G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}\\ \hline\cr G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}\\ G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}\\ G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}\end{array}\right).\end{aligned}$ (15)
Here, for simplicity, $A^{\prime}$ denotes the transpose of $A$. With $\Delta=\{0,1,2,4,5\}$,
$\Gamma=\{(i,j)\in[L/2]^{2}\mid i+j\not\equiv 3\ (\bmod\,6)\}=[L/2]^{2}\setminus\{(0,3),(1,2),(2,1),(3,0),(4,5),(5,4)\}.$ (16)
For the active matrices $H_{X}$ and $H_{Z}$ with $J=3$,
$H_{X}H_{Z}^{\mathsf{T}}=\bigl(\Psi_{(k-i)\bmod\,6}\bigr)_{0\leq i,k\leq 2}=\begin{pmatrix}\Psi_{0}&\Psi_{1}&\Psi_{2}\\ \Psi_{5}&\Psi_{0}&\Psi_{1}\\ \Psi_{4}&\Psi_{5}&\Psi_{0}\end{pmatrix}$.
Thus active orthogonality is equivalent to
$\Psi_{r}=0\qquad(r\in\{0,1,2,4,5\}=\Delta),$
so the $r=3$ interaction can be kept free. The mother-matrix product $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ is
$\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}=\bigl(\Psi_{(k-i)\bmod\,6}\bigr)_{0\leq i,k\leq 5}=\left(\begin{array}[]{lll|lll}\Psi_{0}&\Psi_{1}&\Psi_{2}&\Psi_{3}&\Psi_{4}&\Psi_{5}\\ \Psi_{5}&\Psi_{0}&\Psi_{1}&\Psi_{2}&\Psi_{3}&\Psi_{4}\\ \Psi_{4}&\Psi_{5}&\Psi_{0}&\Psi_{1}&\Psi_{2}&\Psi_{3}\\ \hline\cr\Psi_{3}&\Psi_{4}&\Psi_{5}&\Psi_{0}&\Psi_{1}&\Psi_{2}\\ \Psi_{2}&\Psi_{3}&\Psi_{4}&\Psi_{5}&\Psi_{0}&\Psi_{1}\\ \Psi_{1}&\Psi_{2}&\Psi_{3}&\Psi_{4}&\Psi_{5}&\Psi_{0}\end{array}\right).$
For the $J=3,L=12$ Tanner graph, the block positions
[(0,0),(0,6),(1,6),(1,1),(2,1),(2,7),(1,7),(1,0)][(0,0),(0,6),(1,6),(1,1),(2,1),(2,7),(1,7),(1,0)] (17)
form an 8-cycle when traversed in order. The corresponding cycle word, in the form of Eq. (14), is
$W=F_{0}G_{0}^{-1}G_{5}F_{0}^{-1}F_{5}G_{5}^{-1}G_{0}F_{5}^{-1}.$
6.2 Design goal and commutation table
With $L/2=6$ and $J=3$, we have $\Delta=\{0,1,2,4,5\}$ (so $3\notin\Delta$). The design goal is
$\Psi_{r}=0\ (r\in\Delta),\qquad\Psi_{3}\neq 0.$ (18)
This guarantees active orthogonality while breaking mother orthogonality, thereby creating room to prevent the latent part from becoming orthogonal to the active part.
As a schematic commutation table, suppose all pairs $(F_{i},G_{j})$ commute except $(i,j)=(0,3)$. The table below shades commuting pairs in $\Gamma$ with 1:
[…]
Then
$\Psi_{r}=0\ (r\neq 3),\qquad\Psi_{3}=F_{0}G_{3}+G_{3}F_{0}+F_{1}G_{2}+G_{2}F_{1}.$ (20)
A reader might think it is sufficient to keep only $\Psi_{3}$ nonzero while setting $\Psi_{4}=0$. We initially considered this and designed the code so that only $F_{0},G_{3}$ were non-commuting, yielding $\Psi_{3}\neq 0$ and $\Psi_{4}=0$. However, this produced many elementary trapping sets (ETS) formed by connecting $k$ length-8 cycles, which are the most harmful [27]. Representative ETSs are shown in Fig. 1. When we instead designed the code so that $\Psi_{3}\neq 0$ and $\Psi_{4}\neq 0$, making both $F_{0},G_{3}$ and $F_{1},G_{2}$ non-commuting, we substantially reduced connected ETSs for $k\leq 10$.
6.3 Affine permutation construction and choice of $P$
The affine permutations $f_{i},g_{i}$ constructed by the algorithm are listed in Table 1.
Table 1: Constructed APM parameters ($P=768,J=3,L=12$, girth=8)
$i$ $f_{i}(x)$ $g_{i}(x)$
0 $763x+435$ $289x+496$
1 $679x+69$ $257x+640$
2 $397x+330$ $625x+200$
3 $61x+18$ $41x+524$
4 $697x+612$ $193x+672$
5 $373x+246$ $449x+672$
If $P$ is a prime power, there is no set of permutations $\sigma_{0},\sigma_{1},\tau_{0},\tau_{1}$ on $[P]$ that simultaneously satisfy: $\sigma_{0}$ commutes with $\tau_{0}$ but not with $\tau_{1}$, and $\sigma_{1}$ does not commute with $\tau_{0}$ but commutes with $\tau_{1}$. Using the standard semidirect-product representation of $\mathrm{AGL}_{1}(\mathbb{Z}/p^{k}\mathbb{Z})$, $(a,b):x\mapsto ax+b$, and the quadratic commutation condition, one can show this by analyzing the resulting zero-product conditions with $p$-adic valuation, which contradict non-commutativity. Therefore we choose $P=768=3\times 2^{8}$.
6.4 ETS library construction
In our initial experiments, ETSs of the form shown in Fig. 1 were the dominant cause of BP stalls. Since the constructed code has girth 8, we first enumerate 8-cycles and then build an ETS library by connecting 8-cycles to generate ETSs isomorphic to these patterns. The library is not a collection of only the ETSs encountered in decoding; instead, we enumerate all ETSs isomorphic to the dominant patterns and add them exhaustively.
Figure 1(a) is a $(6,2)$ ETS obtained by connecting two length-8 cycles, and Fig. 1(b) is a $(12,2)$ ETS obtained by connecting five length-8 cycles. Figure 2 shows a $(8,2)$ ETS formed by attaching a length-4 check–variable path between the two odd checks of a $(6,2)$ ETS. For the constructed code with $P=768,J=3,L=12$, the enumerator found $(6,2)$ ETSs: $X=48$, $Z=16$ (total 64); $(12,2)$ ETSs: $X=23$, $Z=0$ (total 23); and $(8,2)$ ETSs of the path4 type: $X=48$, $Z=0$ (total 48).
The parameters $f_{i},g_{i}$ constructed by the algorithm are listed in Table 1 (indices start at 0). Each $f_{i},g_{i}$ is an affine permutation $x\mapsto ax+b$ on $\mathbb{Z}_{P}$ with $\gcd(a,P)=1$. The resulting Tanner graph has girth 8 (no 4- or 6-cycles). We also verified that small $(a,2)$ ETSs obtained by connecting length-8 cycles do not appear up to a prescribed search limit $a\leq A$ (details omitted).
(a) $(6,2)$ ETS (two connected 8-cycles)
(b) $(12,2)$ ETS (five connected 8-cycles)
Figure 1: ETSs formed by connecting $k=2,5$ length-8 cycles. Filled check nodes have degree 1 (unsatisfied). Counts for $P=768,J=3,L=12$: $(6,2)$ ETS (X=48, Z=16) and $(12,2)$ ETS (X=23, Z=0).
Figure 2: Tanner graph of an ETS with (8,2). The internal subgraph matches the $(6,2)$ ETS in Fig. 1(a), and the length-4 path connects its former degree-1 checks (C0:599 and C0:542). Count for $P=768,J=3,L=12$: $(8,2)$ ETS (X=48, Z=0).
6.5 Minimum-distance evaluation
Let $\boldsymbol{u},\boldsymbol{v}\in\mathbb{F}_{2}^{3P}$ be coefficient vectors satisfying $\operatorname{diag}(\Psi_{3}^{\mathsf{T}},\Psi_{3}^{\mathsf{T}},\Psi_{3}^{\mathsf{T}})\boldsymbol{u}=0$ and $\operatorname{diag}(\Psi_{3},\Psi_{3},\Psi_{3})\boldsymbol{v}=0$. Define $\boldsymbol{x}:=\boldsymbol{u}^{\mathsf{T}}\tilde{H}_{X}$ and $\boldsymbol{z}:=\boldsymbol{v}^{\mathsf{T}}\tilde{H}_{Z}$ using the latent matrices $\tilde{H}_{X}$ and $\tilde{H}_{Z}$. Here $H_{X}$ and $H_{Z}$ are the active matrices, so $\boldsymbol{u}\in\mathrm{Ker}\!\bigl(H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\bigr)$ and $\boldsymbol{v}\in\mathrm{Ker}\!\bigl(H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\bigr)$. With $L/2=6,J=3$ and $\Psi_{r}=0\ (r\neq 3)$,
$H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=\operatorname{diag}(\Psi_{3}^{\mathsf{T}},\Psi_{3}^{\mathsf{T}},\Psi_{3}^{\mathsf{T}}),\qquad H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=\operatorname{diag}(\Psi_{3},\Psi_{3},\Psi_{3}).$ (21)
Commutativity makes the terms for $u=2,3,4,5$ vanish, so
$\Psi_{3}=F_{0}G_{3}+G_{3}F_{0}+F_{1}G_{2}+G_{2}F_{1},$ (22)
and hence
$\Psi_{3}^{\mathsf{T}}=F_{0}^{\mathsf{T}}G_{3}^{\mathsf{T}}+G_{3}^{\mathsf{T}}F_{0}^{\mathsf{T}}+F_{1}^{\mathsf{T}}G_{2}^{\mathsf{T}}+G_{2}^{\mathsf{T}}F_{1}^{\mathsf{T}}.$ (23)
Thus $\Psi_{3}$ is a sum of four permutation matrices with disjoint supports, so both row and column weights are 4. Therefore any codeword $\boldsymbol{u}$ of $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}$ decomposes as $\boldsymbol{u}=(\boldsymbol{u}_{3},\boldsymbol{u}_{4},\boldsymbol{u}_{5})$ where each $\boldsymbol{u}_{i}$ is a codeword of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$. From the structure of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ and $\mathrm{Ker}(\Psi_{3})$, the support splits and weights are multiples of 4. The minimum distance of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ is 4. We confirmed the existence of a weight-48 $\boldsymbol{x}$ constructed from a weight-4 codeword of $\Psi_{3}^{\mathsf{T}}$. Therefore $d_{X}^{(\mathrm{res})},d_{Z}^{(\mathrm{res})}\leq 48$, and consequently $d_{\min}\leq 48$. We do not prove a matching lower bound. Randomized searches did not find smaller logical errors, and large-scale decoding experiments did not exhibit logical failures, so we conjecture that this upper bound is tight.
6.6 FER performance
Figure 3 reports the FER of the constructed code under BP with post-processing. For reference, we also include the hashing bound for the depolarizing channel in the plot. At $p=4\%$, the FER reaches $10^{-8}$. For each point we collected at least 50 error events; 95% confidence intervals are plotted but may be visually indistinguishable because they are narrow. For FER above $10^{-7}$, all failures left hundreds of remaining errors, and we did not observe cases where the decoder returned low-weight logical errors (including higher-weight ones), which is common in small-distance codes. Around $10^{-8}$, an error-floor tendency begins to appear, and the dominant failures are stalls trapped by trapping sets of size on the order of tens.
Figure 3: FER curve of the constructed code girth-8, (3,12)-regular $9216,4612,\leq 48]]$, ($P=768,J=3,L=12$) with error bars (95% confidence). Here $n=PL$, the active matrix ranks are $\mathrm{rank}(H_{X})=2302$ and $\mathrm{rank}(H_{Z})=2302$, hence $k=4612$.